The acceleration of a particle starting from rest varies with time as a = αt + β. The velocity of the particle at time t is
 

Grade:11

1 Answers

Aditya Vaidya
2 years ago
In this question, we are given acceleration as a function of time. To find velocity as a function of time we can integrate the function as follows:
\alpha t+\beta….given
a=dv/dt=\alpha t+\beta
\int dv=\int \alpha t+\beta dt
v=\alpha t^2/2+\beta t